# Zero-Shot Character Identification and Speaker Prediction in Comics via Iterative Multimodal Fusion

### v2

@@ Abstract.
*This work was conducted when the first author was a research intern at Mantra Inc.
@@ 1. Introduction
Previous studies on speaker prediction in comics primarily addressed the correspondence between character regions and speech bubble regions in images (Rigaud et al., 2015; Li et al., 2023), but they did not determine the speakers’ names. If we want to know the character names of the speakers, we first need to identify the character names of the character regions. The straightforward supervised approach involves training models to detect and classify characters in images (Ogawa et al., 2018; Zheng et al., 2020). However, this approach requires annotations for each comic with different characters, which is impractical in the fast-paced comic industry that releases thousands of titles annually.
@@ 2.1. Comic Speaker Prediction
Previous studies of comic speaker prediction focus on predicting the correspondence between character bounding boxes and text regions. Conventional methods are rule-based, primarily relying on the distance between the character and the text region (Rigaud et al., 2015). They disregard visual and textual context information, causing them to fail in cases where the speaker is not closest to the text region. Recently, Li et al. employ scene graph generation (SGG) models (Li et al., 2023) known for their effectiveness in visual relationship detection tasks (Johnson et al., 2015; Xu et al., 2017; Zellers et al., 2018; Tang et al., 2020), to predict the correspondence in a more robust way by using visual semantic information such as the face angle. While the previous studies focused only on predicting the correspondence between character bounding boxes and text regions, we propose a new task to predict speaker identities, i.e., character names. This allows more applications such as text-to-speech with character-specific voices. Furthermore, while previous studies have utilized only visual information, we are the first to utilize both visual and textual information.
@@ 2.3. Large Language Models
The recent success of ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) has demonstrated the power of LLMs in understanding, generating, and interpreting human language with remarkable accuracy. Inspired by these advancements, we have pioneered the application of LLMs to the dialogue analysis in comics. Alongside the development of LLMs, large multimodal models (LMMs) such as LLaVA (Liu et al., 2023b, a) and MiniGPT-4 (Zhu et al., 2023) have garnered attention for processing multimodal information. In spite of their capabilities in visual understanding and reasoning, currently, they can only handle a single or a few images as input. Comic analysis requires comprehension across longer contexts, such as multiple pages. Also, it involves learning character identity throughout the book or series. We thus propose an iterative framework to integrate multimodality and longer contexts into LLMs inference for this task.
@@ 3.1. Problem Settings
Data preprocessing. Before initiating our main pipeline, we conducted a series of preprocessing steps. First, character regions $\mathcal{C}$ and text regions $\mathcal{D}$ are obtained using object detectors for general comic character and text category, which are shown to achieve high accuracy (Ogawa et al., 2018). Subsequently, we derived the initial relationship scores $\mathcal{R}$ from $\mathcal{I}$, $\mathcal{C}$, and $\mathcal{D}$ using the SGG models (Li et al., 2023). Relationship scores =$\mathcal{R}=\{r_{i\to j}\}_{i,j}$ represent the correspondences between character and text regions, where $r_{i\to j}\in(0,1)$ is the confidence that the character $c_{i}$ is the speaker of the text $d_{j}$. This is used to propagate labels between the character and text regions. Additionally, we extracted texts $\mathcal{T}=\{t_{j}\}_{j=1}^{M}$ from text regions $\mathcal{D}$ utilizing optical character recognition (OCR). Through text analysis by LLMs, we obtained the list of character names $\mathcal{N}$ from $\mathcal{T}$.
@@ 3.3. Initial Speaker Prediction
We first predict speakers using only the textual content using LLMs as $F_{0}(\mathcal{T,N})$. GPT-4 (OpenAI, 2023) is used in this work. The input and outputs of LLMs are shown in Figure 3. We input the dialogues $\mathcal{T}$ and the list of appearing characters’ names $\mathcal{N}$ into the LLMs. Due to output text length limitations of GPT-4, we split the dialogues into chunks. To compensate for the missing context about the story in each chunk, we first extract context information about the story summary and character profile from $\mathcal{T}$ and $\mathcal{N}$ using LLMs. This context is then fed into LLMs with each chunk as shown in Figure 3. Note that speaker candidates $\mathcal{Y}_{\mathrm{pseudo}}$ are not input in the initial prediction. As the output, we let LLMs output both character IDs and character names to get a stable output. These output character IDs are converted into $\mathcal{Y}$. In addition, we let LLMs output integer confidence scores from 1 to 5, and we exclude data with low confidence from subsequent steps.
@@ 3.4. Multimodal Character Identification
The first step is pseudo label generation with the label propagation module $H_{\mathrm{t\rightarrow c}}$. For each character region, we select the character-text pair that has the highest relationship score among all combinations involving this character region. Then, we use the selected text region’s label as the pseudo label of the character region. To get higher precision, we set a threshold for the confidence of text region’s label obtained by LLMs. In our experiments, removing data with confidence less than 3 lead to higher precision without reducing recall.
In the second step, we construct a character classifier to identify the characters with unknown labels. We train an image classifier $f_{\mathrm{classify}}$ with pseudo labels $\mathcal{X}_{\mathrm{pseudo}}$ and then obtain $\mathcal{X}$ by applying $f_{\mathrm{classify}}$ to character regions $\mathcal{C}$. We use ResNet50 (He et al., 2016) as the image classifier.
@@ 3.5. Multimodal Speaker Prediction
In the multimodal speaker prediction, we first generate pseudo labels for text regions $\mathcal{Y}_{\mathrm{pseudo}}$ from $\mathcal{X}$ by using the label propagation module $H_{\mathrm{c\rightarrow t}}$ in the same step as $H_{\mathrm{t\rightarrow c}}$. We treat these pseudo labels as candidate speakers for the texts, taking the confidence of character identification to be the confidence for the speaker candidates. Confidences with less than 0.5 are filtered out from candidates.
@@ 3.6. Relationship Rescoring
Relationship scores $\mathcal{R}$ are updated based on the predicted labels $\mathcal{X}$ and $\mathcal{Y}$. Since character labels are not considered in the initial relationship scores, more accurate relationship scores can be obtained by considering character labels. We use a simple rescoring method: $r_{i,j}$ is multiplied by scale $s$ if $x_{i}=y_{j}$ and divided by $s$ if $x_{i}\neq y_{j}$. The scale $s$ is computed from confidences of labels $p_{x_{i}}$ and $p_{y_{j}}$ as $\min(1,\lambda p_{x_{i}}p_{y_{j}})$, where $\lambda$ is a hyper-parameter that is set to 2 in our experiments. With this step, pairs that are predicted as having the same character labels gets higher relationship scores, and vice versa. This is similar to the process where humans recognize characters and predict the speaker-text correspondences when reading comics.
@@ 4.1. Experimental Setup
For the character regions, we used the body region annotations (Ogawa et al., 2018), which represent the entire bodies of the characters. We preferred this method because whole-body regions offer more information (such as clothing and body types) for character identification than face regions. Besides, comic datasets often include characters with body-only annotations, such as those shown from the back or non-human characters, which necessitates the use of body regions for identification. For the character labels associated with the text regions, we used Manga109Dialog annotations (Li et al., 2023).
@@
Relationship Prediction. Our method used the relationship scores $\mathcal{R}$ to show correspondences between character and text regions. We compared three types of initial relationship scores: SGG: a deep learning-based relationship prediction using the scene graph generation (SGG) model (Li et al., 2023), Distance: a rule-based method using the distance between the center coordinates of the character and text regions (Rigaud et al., 2015), and GT: ground truth annotations of Manga109Dialog. We used SGG in the main experiments and used Distance and GT in the ablation study. The confidence of SGG model was used as the score for SGG, and the reciprocal of the distance was used as the score for Distance. In the case of GT, we took the relationship score to be 1.0 for all pairs.
@@
Baselines. Since no existing method can predict character labels in a zero-shot setup, we constructed our own baselines. For character identification task, we first group character regions of each comic by the clustering of deep features obtained using the Manga109 pre-trained model explained above. K-means clustering (Lloyd, 1982) with $k$=$\lvert\mathcal{N}\rvert$ is used. Since it is impossible to assign character labels to each cluster without text information, we map each cluster to the character labels using ground truth so that accuracy is maximized, which can be regarded as an upper bound of the clustering approach. The labels of character regions are converted into those of text regions using relationship prediction methods (Rigaud et al., 2015; Li et al., 2023), which are referred to as K-means + Distance and K-means + SGG, respectively in Table 1.
@@ 4.2. Main Results
Across all test sets, the results from iteration 2 showed an improvement over iteration 1. In iteration 3, while the accuracy for Easy data continued to increase, there was a slight decline observed for Hard and Total data. This suggests that there are limits to the accuracy gains in relationship detection from further iterations To investigate this further, we do an analysis using the ground truth relationships in Table 1 (b), where the accuracy of both speaker prediction and character identification keeps increasing with further iterations. These results suggest our iterative process is more effective in the case that the prediction of relationships between text and character regions is accurate.
@@ 4.3. Ablation Study
LLM Prompts. Table 4 shows the results for different LLM prompts in the speaker prediction. The three options in the table correspond to (1) ctx: context information about characters and stories, (2) cand: speaker candidates, and (3) prob: the probability of candidates. Their respective improvements in accuracy are 4.7%, 5.5%, and 2.0%. Each option introduces information not present in the dialogues of each chunk: context provides whole-story information, while the speaker candidates and their probabilities are obtained from the images. These options offer complementary information to each other. The combination of all three options, as shown in Figure 3, performed best.
[…]
Relationship Rescoring. Table 5 shows the effect of the proposed relationship rescoring on relationship prediction and speaker prediction. We evaluated the relationship prediction in the manner described in the previous work (Li et al., 2023): for each text region, the corresponding character region with the highest relationship score was selected. Then, the accuracy was calculated by dividing the number of correctly matched regions by the total number of text regions. Our approach with rescoring achieved an accuracy 79.8%, which is better than previous methods. Moreover, it increased speaker prediction accuracy. In contrast to the previous work that doesn’t consider character labels, our method can predict the relationships better by predicting character labels of both images and texts.
@@
Table 4. Speaker prediction accuracy with different prompts.
@@ 4.4. Zero-shot Evaluation
We evaluated our method in entirely zero-shot settings using only input images. The bounding boxes of the character and the text regions were detected by using the object detection with Faster R-CNN (Ren et al., 2015), and dialogues were extracted by using OCR in the manner described in (Hinami et al., 2021). We extracted a list of character names directly from the dialogues using GPT-4. These names then served as target labels for speaker prediction and character identification tasks.
Table 6 shows the experimental results under entirely zero-shot settings. In assessing the experimental results, we manually created mappings between the extracted names and the true names. Each predicted region is counted as correct if it was detected with and IoU >0.5 and was correctly labeled. Our method achieved 38.7% and 35.0% in accuracy for speaker prediction and character identification tasks, respectively. This is lower than the results shown in Table 1 due to errors in object detection, OCR, and character name extraction. Particularly, the character name extraction significantly affects performance because characters whose names are not extracted cannot be predicted. To show the effect of character name extraction on the performance, we compute the upper bound in Table 6, which is the accuracy under ideal conditions, where all labels of extracted names are perfectly predicted. The recall for extracted names was 62.7%, thereby limiting the upper bounds to 67.3% for speaker prediction and 63.9% for character identification. It should be noted that there are several titles in which character names are rarely mentioned in dialogues. In that case, it is difficult even for humans to predict the names, which poses the challenges in entirely zero-shot settings. On the other hand, we achieve promising results in titles where character names appear frequently in dialogues and the relationships between characters and texts are clear. Figure 1 shows the results of our approach in entirely zero-shot settings from one specific book, which demonstrate promising first step of zero-shot character identification and speaker prediction of unseen comics.
@@
Table 6. Accuracy in an entirely zero-shot setting.

### v4

@@ Abstract.
*This work was conducted when the first author was a research intern at Mantra Inc. For more details, please visit our project page: https://liyingxuan1012.github.io/zeroshot-speaker-prediction.
@@ 1. Introduction
Previous studies on speaker prediction in comics primarily addressed the correspondence between character regions and speech bubble regions in images (Rigaud et al., 2015; Li et al., 2024), but they did not determine the speakers’ names. If we want to know the character names of the speakers, we first need to identify the character names of the character regions. The straightforward supervised approach involves training models to detect and classify characters in images (Ogawa et al., 2018; Zheng et al., 2020). However, this approach requires annotations for each comic with different characters, which is impractical in the fast-paced comic industry that releases thousands of titles annually.
@@ 2.1. Comic Speaker Prediction
Previous studies of comic speaker prediction focus on predicting the correspondence between character bounding boxes and text regions. Conventional methods are rule-based, primarily relying on the distance between the character and the text region (Rigaud et al., 2015). They disregard visual and textual context information, causing them to fail in cases where the speaker is not closest to the text region. Recently, Li et al. employ scene graph generation (SGG) models (Li et al., 2024) known for their effectiveness in visual relationship detection tasks (Johnson et al., 2015; Xu et al., 2017; Zellers et al., 2018; Tang et al., 2020), to predict the correspondence in a more robust way by using visual semantic information such as the face angle. While the previous studies focused only on predicting the correspondence between character bounding boxes and text regions, we propose a new task to predict speaker identities, i.e., character names. This allows more applications such as text-to-speech with character-specific voices. Furthermore, while previous studies have utilized only visual information, we are the first to utilize both visual and textual information.
@@ 2.3. Large Language Models
The recent success of ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) has demonstrated the power of LLMs in understanding, generating, and interpreting human language with remarkable accuracy. Inspired by these advancements, we have pioneered the application of LLMs to the dialogue analysis in comics. Alongside the development of LLMs, large multimodal models (LMMs) such as LLaVA (Liu et al., 2023b, a) and MiniGPT-4 (Zhu et al., 2023) have garnered attention for processing multimodal information. Although these models show strong performance on visual understanding and reasoning within a single panel, they struggle with understanding content across multiple panels (Ikuta et al., 2024). Comic analysis requires comprehension across longer contexts, such as multiple pages. Also, it involves learning character identity throughout the book or series. We thus propose an iterative framework to integrate multimodality and longer contexts into LLMs inference for this task.
@@ 3.1. Problem Settings
Data preprocessing. Before initiating our main pipeline, we conducted a series of preprocessing steps. First, character regions $\mathcal{C}$ and text regions $\mathcal{D}$ are obtained using object detectors for general comic character and text category, which are shown to achieve high accuracy (Ogawa et al., 2018). Subsequently, we derived the initial relationship scores $\mathcal{R}$ from $\mathcal{I}$, $\mathcal{C}$, and $\mathcal{D}$ using the SGG models (Li et al., 2024). Relationship scores =$\mathcal{R}=\{r_{i\to j}\}_{i,j}$ represent the correspondences between character and text regions, where $r_{i\to j}\in(0,1)$ is the confidence that the character $c_{i}$ is the speaker of the text $d_{j}$. This is used to propagate labels between the character and text regions. Additionally, we extracted texts $\mathcal{T}=\{t_{j}\}_{j=1}^{M}$ from text regions $\mathcal{D}$ utilizing optical character recognition (OCR). Through text analysis by LLMs, we obtained the list of character names $\mathcal{N}$ from $\mathcal{T}$.
@@ 3.3. Initial Speaker Prediction
We first use GPT-4 (OpenAI, 2023) to predict speakers only based on textual content, denoted as $F_{0}(\mathcal{T,N})$. The input and outputs of LLMs are shown in Figure 3. We input the dialogues $\mathcal{T}$ and the list of appearing characters’ names $\mathcal{N}$ into the LLMs. Due to output text length limitations of GPT-4, we split the dialogues into chunks. To compensate for the missing context about the story in each chunk, we first extract context information about the story summary and character profile from $\mathcal{T}$ and $\mathcal{N}$ using LLMs. This context is then fed into LLMs with each chunk as shown in Figure 3. Note that speaker candidates $\mathcal{Y}_{\mathrm{pseudo}}$ are not input in the initial prediction. As the output, we let LLMs output both character IDs and character names to get a stable output. These output character IDs are converted into $\mathcal{Y}$. In addition, we let LLMs output integer confidence scores from 1 to 5, and we exclude data with low confidence from subsequent steps.
@@ 3.4. Multimodal Character Identification
The first step is pseudo label generation with the label propagation module $H_{\mathrm{t\rightarrow c}}$. For each character region, we select the character-text pair that has the highest relationship score among all combinations involving this character region. Then, we use the selected text region’s label as the pseudo label of the character region. To get higher precision, we set a threshold for the confidence of text region’s label obtained by LLMs. In our experiments, removing data with confidence less than 3 lead to higher precision without reducing recall. In the second step, we construct a character classifier to identify the characters with unknown labels. We train an image classifier $f_{\mathrm{classify}}$ with pseudo labels $\mathcal{X}_{\mathrm{pseudo}}$ and then obtain $\mathcal{X}$ by applying $f_{\mathrm{classify}}$ to character regions $\mathcal{C}$. We use ResNet50 (He et al., 2016) as the image classifier.
@@ 3.5. Multimodal Speaker Prediction
In the multimodal speaker prediction, we first generate pseudo labels for text regions $\mathcal{Y}_{\mathrm{pseudo}}$ from $\mathcal{X}$ by using the label propagation module $H_{\mathrm{c\rightarrow t}}$ in the same step as $H_{\mathrm{t\rightarrow c}}$. These pseudo labels serve as candidate speakers for the texts, with character identification confidence serving as the confidence for these candidates. Confidences with less than 0.5 are filtered out from candidates.
@@ 3.6. Relationship Rescoring
Relationship scores $\mathcal{R}$ are updated based on the predicted labels $\mathcal{X}$ and $\mathcal{Y}$. Since character labels are not considered in the initial relationship scores, more accurate relationship scores can be obtained by considering character labels. We use a simple rescoring method: $r_{i,j}$ is multiplied by scale $s$ if $x_{i}=y_{j}$ and divided by $s$ if $x_{i}\neq y_{j}$. The scale $s$ is computed from confidences of labels $p_{x_{i}}$ and $p_{y_{j}}$ as $\min(1,\lambda p_{x_{i}}p_{y_{j}})$, where $\lambda$ is a hyper-parameter that is set to 2 in our experiments. With this step, pairs with the same character labels get higher relationship scores, and vice versa. This is similar to the process where humans recognize characters and predict the speaker-text correspondences when reading comics.
@@ 4.1. Experimental Setup
For the character regions, we used the body region annotations (Ogawa et al., 2018), which represent the entire bodies of the characters. We preferred this method because whole-body regions offer more information (such as clothing and body types) for character identification than face regions. Besides, comic datasets often include characters with body-only annotations, such as those shown from the back or non-human characters, which necessitates the use of body regions for identification. For the character labels associated with the text regions, we used Manga109Dialog annotations (Li et al., 2024).
@@
Relationship Prediction. Our method used the relationship scores $\mathcal{R}$ to show correspondences between character and text regions. We compared three types of initial relationship scores: SGG: a deep learning-based relationship prediction using the scene graph generation (SGG) model (Li et al., 2024), Distance: a rule-based method using the distance between the center coordinates of the character and text regions (Rigaud et al., 2015), and GT: ground truth annotations of Manga109Dialog. We used SGG in the main experiments and used Distance and GT in the ablation study. The confidence of SGG model was used as the score for SGG, and the reciprocal of the distance was used as the score for Distance. In the case of GT, we took the relationship score to be 1.0 for all pairs.
@@
Baselines. Since no existing method can predict character labels in a zero-shot setup, we constructed our own baselines. For character identification task, we first group character regions of each comic by the clustering of deep features obtained using the Manga109 pre-trained model explained above. K-means clustering (Lloyd, 1982) with $k$=$\lvert\mathcal{N}\rvert$ is used. Since it is impossible to assign character labels to each cluster without text information, we map each cluster to the character labels using ground truth so that accuracy is maximized, which can be regarded as an upper bound of the clustering approach. The labels of character regions are converted into those of text regions using relationship prediction methods (Rigaud et al., 2015; Li et al., 2024), which are referred to as K-means + Distance and K-means + SGG, respectively in Table 1.
@@ 4.2. Main Results
Across all test sets, the results from iteration 2 showed an improvement over iteration 1. In iteration 3, while the accuracy for Easy data continued to increase, there was a slight decline observed for Hard and Total data. This suggests that there are limits to the accuracy gains in relationship detection from further iterations. To investigate this further, we do an analysis using the ground truth relationships in Table 1 (b), where the accuracy of both speaker prediction and character identification keeps increasing with further iterations. These results suggest our iterative process is more effective in the case that the prediction of relationships between text and character regions is accurate.
@@ 4.3. Ablation Study
LLM Prompts.
Table 4. Speaker prediction accuracy with different prompts.
@@
Table 4 shows the results for different LLM prompts in the speaker prediction. The three options in the table correspond to (1) ctx: context information about characters and stories, (2) cand: speaker candidates, and (3) prob: the probability of candidates. Their respective improvements in accuracy are 4.7%, 5.5%, and 2.0%. Each option introduces information not present in the dialogues of each chunk: context provides whole-story information, while the speaker candidates and their probabilities are obtained from the images. These options offer complementary information to each other. The combination of all three options, as shown in Figure 3, performed best.
[…]
Relationship Rescoring. Table 5 shows the effect of the proposed relationship rescoring on relationship prediction and speaker prediction. We evaluated the relationship prediction in the manner described in the previous work (Li et al., 2024): for each text region, the corresponding character region with the highest relationship score was selected. Then, the accuracy was calculated by dividing the number of correctly matched regions by the total number of text regions. Our approach with rescoring achieved an accuracy 79.8%, which is better than previous methods. Moreover, it increased speaker prediction accuracy. In contrast to the previous work that doesn’t consider character labels, our method can predict the relationships better by predicting character labels of both images and texts.
@@ 4.4. Zero-shot Evaluation
We evaluated our method in entirely zero-shot settings using only input images. The bounding boxes of the character and the text regions were detected with Faster R-CNN (Ren et al., 2015), and dialogues were extracted by using OCR in the manner described in (Hinami et al., 2021). We extracted a list of character names directly from the dialogues using GPT-4. These names then served as target labels for speaker prediction and character identification tasks.
Table 6 shows the experimental results under entirely zero-shot settings. In assessing the experimental results, we manually mapped extracted names to true names. Each predicted region is counted as correct if it was detected with an IoU >0.5 and was correctly labeled. Our method achieved 38.7% and 35.0% in accuracy for speaker prediction and character identification tasks, respectively. This is lower than the results shown in Table 1 due to errors in object detection, OCR, and character name extraction. Our experiments demonstrated that the impact of errors from object detection and OCR on overall performance is limited. When provided only with images and a list of character names, we achieved 45.9% and 40.0% in accuracy, respectively. To show the effect of character name extraction on the performance, we computed the upper bound in Table 6, which is the accuracy under ideal conditions, where all labels of extracted names are perfectly predicted. The recall for extracted names was 62.7%, thereby limiting the upper bounds to 67.3% for speaker prediction and 63.9% for character identification. We analyzed the accuracy across different comic titles in an entirely zero-shot setting and found significant variations. These variations arise because character names are rarely mentioned in the dialogues of several titles, posing a challenge that is difficult even for humans. However, we observed promising results in titles where character names appear frequently in dialogues and the relationships between characters and texts are clear, as examples shown in Figure 1.
Table 6. Accuracy in an entirely zero-shot setting.